Explain with examples:
         (i) Atomic number         (ii) Mass number,
         (iii) Isotopes and             (iv) Isobars.
ik;.iyluuyg;p8yip[;9i[;puipy8ufply8u

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student



(i) Atomic number:The atomic number of an element is equal to the number of protons in the nucleus of its atom.
e.g., Oxygen has 6 protons hence atomic no. = 6.
(ii) Maas number:The mass number of an atom is equal to the number of protons and neutrons in its nucleus.
Nucleons= number of protons + number of neutrons
Example: Protons + Neutrons = Nucleus = Mass number
6 + 6 = 12
(iii)Isotopes:Isotopes are atoms of the same element which have different mass number but same atomic number.

(iv)Isobars:Isobars are atoms having the same mass number but riifteren atomic numbers.

Both calcium and argon have same mass number but different storm number.
Two uses of isotopes are:
(i) An isotope of iodine is used in the treatment of goitre.
(ii) An isotope of uranium is used as a fuel in nuclear reactors.


Thanks